# Meta-rater: A Multi-dimensional Data Selection Method for Pre-training Language Models

### v1

@@ Abstract
The composition of pre-training datasets for large language models (LLMs) remains largely undisclosed, hindering transparency and efforts to optimize data quality—a critical driver of model performance. Current data selection methods, such as natural language quality assessments, diversity-based filters, and classifier-based approaches, are limited by single-dimensional evaluation or redundancy-focused strategies. To address these gaps, we propose PRRC to evaluate data quality across Professionalism, Readability, Reasoning, and Cleanliness. We further introduce Meta-rater, a multi-dimensional data selection method that integrates these dimensions with existing quality metrics through learned optimal weightings. Meta-rater employs proxy models to train a regression model that predicts validation loss, enabling the identification of optimal combinations of quality scores. Experiments demonstrate that Meta-rater doubles convergence speed for 1.3B parameter models and improves downstream task performance by 3.23%, with scalable benefits observed in 3.3B models trained on 100B tokens. Additionally, we release the annotated SlimPajama-627B dataset, labeled across 25 quality metrics (including PRRC), to advance research in data-centric LLM development. Our work establishes that holistic, multi-dimensional quality integration significantly outperforms conventional single-dimension approaches, offering a scalable paradigm for enhancing pre-training efficiency and model capability.
@@
Xinlin Zhuang2,1thanks: These authors contributed equally to this work., Jiahui Peng1∗, Ren Ma1∗thanks: Project lead., Yinfan Wang1, Tianyi Bai1, Xingjian Wei1, Jiantao Qiu1, Chi Zhang1, Ying Qian2, Conghui He1thanks: Corresponding author. 1 Shanghai Artificial Intelligence Laboratory 2 School of Computer Science and Technology, East China Normal University xinlinzhuang@stu.ecnu.edu.cn, heconghui@pjlab.org.cn
@@ 1 Introduction
Existing pre-training data selection methods can be categorized into three primary approaches: natural language quality-based methods Rae et al. (2021); Weber et al. (2024); Xie et al. (2023b); Ankner et al. (2024), diversity-based methods Abbas et al. (2023); He et al. (2024b); Zhang et al. (2025), and classifier-based methods Wettig et al. (2024); Penedo et al. (2024). Alternative strategies such as MATES Yu et al. (2024) utilize influence scores Park et al. (2023) and Rho Lin et al. (2024) formulates data selection at the token level. However, these methods exhibit inherent constraints - natural language quality assessment focuses on superficial text characteristics, diversity-based approaches prioritize redundancy reduction over intrinsic quality evaluation, and classifier-based techniques operate through single-dimensional quality filters. This raises the fundamental question: How can we systematically integrate complementary quality dimensions to achieve holistic data selection?
To address this gap, we develop four novel evaluation dimensions PRRC (Professionalism, Readability, Reasoning, and Cleanliness) to expand current quality metrics. Utilizing these dimensions, we introduce Meta-rater, a model-based framework that strategically integrates multiple quality scores for optimal data selection. Meta-rater operates by training small proxy models and fitting a model on their data, thereby deriving optimal combination of various quality scores. Empirical validation demonstrates Meta-rater’s efficacy across model scales: for 1.3B parameter models trained on 30B tokens, it achieves twice the convergence speed compared to random selection and a 3.23% overall performance improvement. Scalability is evidenced with 3.3B models trained on 100B tokens, where performance gains persist. These results substantiate that integrating multi-dimensional quality metrics surpasses conventional single-dimension approaches, establishing a new paradigm for data curation in LLM development.
@@
  • Empirical Validation: We demonstrate Meta-rater’s practical impact—doubled convergence speed and 3.23% downstream task improvement for 1.3B models—with scalability validated on 3.3B models trained on 100B tokens.
@@ 2 Related Work
For natural language quality-based methods, Gopher Rae et al. (2021) and RedPajama Weber et al. (2024) propose empirical rules like controlling the ratio of word and number token in texts to improve language modeling. Additionally, previous works  Muennighoff et al. (2024); Wenzek et al. (2020) have shown that selecting data with perplexity (PPL) scores on validation datasets can lead to superior performance on downstream tasks compared to using the entire dataset. Another notable method is DSIR Xie et al. (2023b), which streamlines the selection process by employing hashed N-gram features (named as data importance scores) to efficiently identify high-quality data within large datasets. Meanwhile, another line of works utilize clustering Zhang et al. (2025) or deduplication Abbas et al. (2023); He et al. (2024b) to enhance diversity of pre-training datasets.
@@ 3.1 Task Formulation
Previously, this task was typically completed by top-k selection based on a single quality score measuring one dimension. In this work, we extend this task to incorporate multiple quality scores covering different dimensions. The challenge then becomes how to aggregate these scores to derive an final data quality score:
@@ 3.2 Meta-rater Design
    Generate a set of random weights Generate a set of random weights $\mathbf{w}_{i}=\{w_{i1},w_{i2},\dots,w_{im}\}$ for for $m$ quality scores, where each weight represents the importance of a specific quality dimension. quality scores, where each weight represents the importance of a specific quality dimension.
@@ 3.3 Data Quality Scores
    Reasoning. With the introduction of OpenAI’s o1 model [^1] https://openai.com/index/introducing-openai-o1-preview/, LLMs have transitioned into the era of reasoning models. Research by DeepSeek has shown that smaller language models can achieve reasoning capabilities on par with LLMs that are ten times their size by leveraging supervised fine-tuning on high-reasoning data Guo et al. (2025). This finding highlights the critical importance of data rich in reasoning complexity. To address this, we developed this dimension to identify data exhibiting exceptional reasoning depth. Such data typically involves multi-step logical reasoning, thorough analysis, and requires readers to synthesize diverse information to form well-rounded conclusions.
@@ 4.1 Experimental Setup
Table 1: Performance of data selection methods on downstream tasks. For Meta-rater, the number in parentheses ( ) indicates the number of quality scores used. We report performance improvements relative to random sampling of 30B tokens, with the best result highlighted and the second best result underlined in each column. Model refers to model-based ratings, while All denotes the inclusion of all quality scores.
@@
To comprehensively assess the capabilities of pre-trained models, we conduct holistic evaluations on various downstream tasks covering three significant categories:
  • General Knowledge: ARC-Challenge Clark et al. (2018), ARC-Easy, and SciQ Welbl et al. (2017).
  • Commonsense Reasoning: HellaSwag Zellers et al. (2019), SIQA Sap et al. (2019), and WinoGrande Sakaguchi et al. (2020).
  • Reading Comprehension: RACE Lai et al. (2017) and OpenbookQA Mihaylov et al. (2018).
Evaluations are conducted using the lm-evaluation-harness Gao et al. (2023) framework with in-context learning setting, and average accuracy is reported for convenient comparison. Further details of evaluation are shown in Appendix E.
@@ 4.2 Results and Analysis
We evaluate all baseline models and those trained using Meta-rater. Evaluation results are presented in Table 1 and full results are provided in Appendix H. Meta-rater achieves the highest performance compared to previous data selection methods. Specifically, it surpasses the Random-30B by a margin of 3.23 (representing a $\sim$ 7.4% relative improvement) in average accuracy and exceeds the previous SOTA method, QuRating Educational Value, by 0.85 (a $\sim$ 1.8% relative gain). Notably, Meta-rater excels across all task categories, highlighting its robustness and versatility in addressing a wide range of downstream tasks.
@@
Meta-rater is computational efficient.
As shown in Figure 1, Meta-rater matches the performance to Random-30B model using only 15B tokens. When consuming 30B tokens, it surpasses the Random-60B model by a margin of 1.31, despite using half the number of tokens. To quantify computational efficiency, we analyze the FLOPs required for quality score rating, Meta-rater construction, and language model pre-training, with detailed breakdowns provided in Appendix F. As shown in Table 2, the FLOPs for Meta-rater constitute only 0.7% of those required to pre-train a 1.3B model. Although the FLOPs for quality score rating are approximately 1.4 times higher than pre-training a 1.3B model, the annotated labels generated are reusable for various purposes and represent a valuable resource for the broader research community. Furthermore, the cost-effectiveness of the rating process becomes increasingly pronounced at larger pre-training scales: it accounts for only 17% of the FLOPs required to pre-train a 3.3B model on 100B tokens. In summary, Meta-rater demonstrates significant advantages in enabling efficient and scalable pre-training.
Scalability with respect to the number of quality scores.
@@
We conduct additional experiments to evaluate Meta-rater’s effectiveness in pre-training a 3.3B model using 100B tokens and a 7.2B model using 150B tokens from scratch, comparing random sampling against Meta-rater with all 25 raters. As shown in Table 3, Meta-rater consistently outperforms random sampling across all three task categories, with the most significant improvement observed in General Knowledge tasks (67.51 vs 64.22). Specifically, when pre-trained on 100B tokens, Meta-rater achieves an average score of 54.71, surpassing the random sampling baseline of 52.98 by 1.73 ($\sim$ 3.3% relative improvement). Similarly in Table 4, Meta-rater outperforms random baseline by 3.12 ($\sim$ 6.0% relative improvement) for pre-training 7.2B models. These results underscore Meta-rater’s ability to select higher-quality training data, particularly for scaling to larger models and datasets.
Method G.K. C.R. R.C. Avg.
Random
   50B Tokens 61.46 49.95 32.90 50.00
   100B Tokens 64.22 53.55 35.28 52.98
Meta-rater
   50B Tokens 65.08 50.00 32.97 51.40
   100B Tokens 67.51 54.35 36.06 54.71
Table 3: Performance of 3.3B models with random sampling and Meta-rater on downstream tasks. Abbreviations: G.K. = General Knowledge, C.R. = Commonsense Reasoning, R.C. = Reading Comprehension.
Method G.K. C.R. R.C. Avg.
Random
   50B Tokens 62.12 50.76 33.18 50.63
   150B Tokens 65.10 52.01 35.87 52.12
Meta-rater
   50B Tokens 64.75 51.56 34.00 52.88
   150B Tokens 67.97 54.58 37.14 55.24
Table 4: Performance of 7.2B models with random sampling and Meta-rater on downstream tasks.
@@ 5 Analysis
Table 5: Downstream tasks results comparison of naive rater combination methods.
5.1 Effect of the Number of Proxy Models
@@
Figure 2: Average downstream task performance of Meta-rater with different numbers of proxy models $N$.
@@ 5.2 Effect of Combining Strategies for Quality Scores
We explore two straightforward methods for combining quality scores as alternatives to the Meta-rater: Mean: This approach uses the arithmetic mean of all quality scores, giving equal weight to each score, and Intersection [^3] Due to strict selection criteria, sufficient data for pre-training could only be obtained with 4 QuRating raters and 4 PRRC raters.: This method selects data that meet the criteria for all quality scores.
While both methods provide ways to combine quality scores, they result in only slight performance improvements. The proposed Meta-rater outperforms the Mean method, achieving an average score of 47.01 compared to 44.65 with 25 raters. A similar gap is observed between Meta-rater and the Intersection method with PRRC raters (46.35 vs 45.17). We believe this superiority is due to key limitations of these simple combinations. Specifically, the Mean approach assumes equal importance among all raters, which can lead to suboptimal results when there are imbalances or inconsistencies in scoring. As shown in Appendix G, individual quality score distributions vary significantly, making uniform weighting ineffective. The Intersection method results in excessive data elimination due to strict filtering criteria. A single low score can lead to data exclusion, even if other raters provide high scores. In contrast, the Meta-rater’s weighted aggregation dynamically adjusts rater contributions while preserving data integrity. This adaptive weighting reduces the impact of inconsistent raters and ensures valuable information is retained, leading to more appropriate selection.
@@ 6 Conclusion
This paper highlights the critical importance of pre-training data quality in determining the performance of LLMs. We present Meta-rater, a novel data selection framework that integrates diverse quality scores across multiple dimensions to identify optimal training data. Through extensive experiments, we demonstrate that Meta-rater consistently outperforms existing data selection methods, achieving significant performance gains across various downstream tasks. These results underscore the effectiveness of Meta-rater in enhancing the efficiency and effectiveness of LLM pre-training.

### v3

@@ Abstract
The composition of pre-training datasets for large language models (LLMs) remains largely undisclosed, hindering transparency and efforts to optimize data quality—a critical driver of model performance. Current data selection methods, such as natural language quality assessments, diversity-based filters, and classifier-based approaches, are limited by single-dimensional evaluation or redundancy-focused strategies. To address these gaps, we propose four dimensions to evaluate data quality: professionalism, readability, reasoning, and cleanliness. We further introduce Meta-rater, a multi-dimensional data selection method that integrates these dimensions with existing quality metrics through learned optimal weightings. Meta-rater employs proxy models to train a regression model that predicts validation loss, enabling the identification of optimal combinations of quality scores. Experiments demonstrate that Meta-rater doubles convergence speed for 1.3B parameter models and improves downstream task performance by 3.23%, with advantages that scale to models as large as 7.2B parameters. Our work establishes that holistic, multi-dimensional quality integration significantly outperforms conventional single-dimension approaches, offering a scalable paradigm for enhancing pre-training efficiency and model capability. To advance future research, we release scripts, data, and models at https://github.com/opendatalab/Meta-rater.
@@
Xinlin Zhuang2,1thanks: These authors contributed equally to this work., Jiahui Peng1∗, Ren Ma1∗thanks: Project lead., Yinfan Wang1, Tianyi Bai1, Xingjian Wei1, Jiantao Qiu1, Chi Zhang1, Ying Qian2, Conghui He1thanks: Corresponding author. 1 Shanghai Artificial Intelligence Laboratory 2 School of Computer Science and Technology, East China Normal University xinlinzhuang@stu.ecnu.edu.cn, maren@pjlab.org.cn, heconghui@pjlab.org.cn
@@ 1 Introduction
Existing pre-training data selection methods can be categorized into three primary approaches: natural language quality-based methods Rae et al. (2021); Weber et al. (2024); Xie et al. (2023b); Ankner et al. (2024), diversity-based methods Abbas et al. (2023); He et al. (2024b); Zhang et al. (2025); Bai et al. (2024), and classifier-based methods Wettig et al. (2024); Penedo et al. (2024). Alternative strategies such as MATES Yu et al. (2024) utilize influence scores Park et al. (2023) and Rho Lin et al. (2024) formulates data selection at the token level. However, these methods exhibit inherent constraints - natural language quality assessment focuses on superficial text characteristics, diversity-based approaches prioritize redundancy reduction over intrinsic quality evaluation, and classifier-based techniques operate through single-dimensional quality filters. This raises the fundamental question: How can we systematically integrate complementary quality dimensions to achieve holistic data selection?
To address this gap, we develop four novel evaluation dimensions PRRC (Professionalism, Readability, Reasoning, and Cleanliness) to expand current quality metrics. Utilizing these dimensions, we introduce Meta-rater, a model-based framework that strategically integrates multiple quality scores for optimal data selection. Meta-rater operates by training small proxy models and fitting a model on their data, thereby deriving optimal combination of various quality scores. Empirical validation demonstrates Meta-rater’s efficacy across model scales: for 1.3B parameter models trained on 30B tokens, it achieves twice the convergence speed compared to random selection and a 3.23% overall performance improvement. Scalability is evidenced with 3.3B and 7.2B models, where performance gains persist. These results substantiate that integrating multi-dimensional quality metrics surpasses conventional single-dimension approaches, establishing a new paradigm for data curation in LLM development.
@@
  • Empirical Validation: We demonstrate Meta-rater’s practical impact—doubled convergence speed and 3.23% downstream task improvement for 1.3B models—with scalability validated on 3.3B and 7.2B models.
@@ 2 Related Work
For natural language quality-based methods, Gopher Rae et al. (2021) and RedPajama Weber et al. (2024) propose empirical rules like controlling the ratio of word and number tokens in texts to improve language modeling. Additionally, previous works  Muennighoff et al. (2024); Wenzek et al. (2020) have shown that selecting data with perplexity (PPL) scores on validation datasets can lead to superior performance on downstream tasks compared to using the entire dataset. Another notable method is DSIR Xie et al. (2023b), which streamlines the selection process by employing hashed N-gram features (named as data importance scores) to efficiently identify high-quality data within large datasets. Meanwhile, another line of works utilize clustering Zhang et al. (2025) or deduplication Abbas et al. (2023); He et al. (2024b) to enhance diversity of pre-training datasets.
@@ 3.1 Task Formulation
Previously, this task was typically completed by top-k selection based on a single quality score measuring one dimension. In this work, we extend this task to incorporate multiple quality scores covering different dimensions. The challenge then becomes how to aggregate these scores to derive a final data quality score:
@@ 3.2 Meta-rater Design
    Generate a set of random weights $\mathbf{w}_{i}=\{w_{i1},w_{i2},\dots,w_{im}\}$ for $m$ quality scores, where each weight represents the importance of a specific quality dimension.
@@ 3.3 Data Quality Scores
    Reasoning. With the introduction of OpenAI’s o1 model, LLMs have transitioned into the era of reasoning models. Research by DeepSeek has shown that smaller language models can achieve reasoning capabilities on par with LLMs that are ten times their size by leveraging supervised fine-tuning on high-reasoning data Guo et al. (2025). This finding highlights the critical importance of data rich in reasoning complexity. To address this, we developed this dimension to identify data exhibiting exceptional reasoning depth. Such data typically involves multi-step logical reasoning, thorough analysis, and requires readers to synthesize diverse information to form well-rounded conclusions.
@@ 4.1 Experimental Setup
Table 1: Performance of data selection methods on downstream tasks. For Meta-rater, the number in parentheses ( ) indicates the number of quality scores used. We report performance improvements compared to random sampling of 30B tokens, with the best result highlighted and the second best result underlined in each column. Model refers to model-based ratings, while All denotes the inclusion of all 25 quality scores. Full evaluation results are provided in Appendix H.
@@
To comprehensively assess the capabilities of pre-trained models, we conduct holistic evaluations on various downstream tasks covering three significant categories: General Knowledge (including ARC-Challenge Clark et al. (2018), ARC-Easy, and SciQ Welbl et al. (2017)), Commonsense Reasoning (including HellaSwag Zellers et al. (2019), SIQA Sap et al. (2019), and WinoGrande Sakaguchi et al. (2020)), and Reading Comprehension (including RACE Lai et al. (2017) and OpenbookQA Mihaylov et al. (2018)). Evaluations are conducted using the lm-evaluation-harness Gao et al. (2023) framework with in-context learning setting, and average accuracy is reported for convenient comparison. Further details of evaluation are shown in Appendix E.
@@ 4.2 Results and Analysis
4.2.1 Analysis of Quality Metrics
Analysis of quality score weight distribution.
Table 10 presents the learned weights of all 25 quality scores, revealing significant patterns in how different quality dimensions contribute to model performance. Our findings show that Educational Value emerges as the most influential metric (5.64%), confirming observations from QuRating and FineWeb-Edu research. In contrast, Writing Style has minimal impact (0.05%), which aligns with QuRating’s observation that high Writing Style content failed to outperform random sampling. Despite their simplicity, natural language signals prove valuable, especially those that identify non-alphabetical content. Among our PRRC metrics, Reasoning (4.44%) and Professionalism (4.05%) make substantial contributions, while Cleanliness shows comparatively lower influence (1.17%). These weight distributions demonstrate Meta-rater’s effectiveness in identifying and appropriately weighting quality dimensions according to their genuine impact on downstream performance.
Correlations between quality metrics.
We also analyzed the relationships between quality metrics by calculating Spearman correlation coefficients across all 25 quality scores using 200k examples from SlimPajama, with results visualized in Figure 4. Our analysis reveals three key patterns. First, model-based metrics (including our PRRC) exhibit relatively weak correlation (<0.6) with most existing metrics, indicating they capture distinct aspects of data quality. Second, Natural Language Quality Signals demonstrate strong inter-correlation among features like word count, entropy, and sentence count (>0.85). Third, Data Importance Scores (DSIR) show remarkably high correlation with each other (>0.95) while maintaining low correlation with model-based ratings. These observations highlight that our PRRC metrics and other model-based ratings contribute novel information beyond what traditional statistical features capture, supporting their integration into our comprehensive quality assessment framework.
4.2.2 Results of Pre-trained Models
@@
We evaluate all baseline models and those trained using Meta-rater. Evaluation results are presented in Table 1. Meta-rater achieves the highest performance compared to previous data selection methods. Specifically, it surpasses the Random-30B by a margin of 3.23 in average accuracy and exceeds the previous SOTA method, QuRating Educational Value, by 0.85. Notably, Meta-rater excels across all task categories, highlighting its robustness and versatility in addressing a wide range of downstream tasks. Additionally, we conduct evaluations on knowledge-intensive benchmarks such as MMLU Hendrycks et al. (2021) and NaturalQuestions Kwiatkowski et al. (2019). The results, detailed in Appendix I, align with our primary findings and further confirm Meta-rater’s effectiveness compared to all baseline methods.
@@
Meta-rater is computationally efficient.
As shown in Figure 1, Meta-rater matches the performance of Random-30B model using only 15B tokens. When consuming 30B tokens, it surpasses the Random-60B model by a margin of 1.31, despite using half the number of tokens. To quantify computational efficiency, we analyze the FLOPs required for quality score rating, Meta-rater construction, and language model pre-training, with detailed breakdowns provided in Appendix F. As shown in Table 2, the FLOPs for Meta-rater constitute only 0.7% of those required to pre-train a 1.3B model. Although the FLOPs for quality score rating are approximately 1.4 times higher than pre-training a 1.3B model, the annotated labels generated are reusable for various purposes and represent a valuable resource for the broader research community. Furthermore, the cost-effectiveness of the rating process becomes increasingly pronounced at larger pre-training scales: it accounts for only 17% of the FLOPs required to pre-train a 3.3B model on 100B tokens. In summary, Meta-rater demonstrates significant advantages in enabling efficient and scalable pre-training.
Scalability on the number of quality scores.
@@
We conduct additional experiments to evaluate Meta-rater’s effectiveness when scaling to larger models and datasets. Specifically, we pre-train 3.3B models using 100B tokens and 7.2B model using 150B tokens from scratch, comparing random sampling against Meta-rater with all 25 raters. As shown in Table 3, Meta-rater consistently outperforms random sampling across all model sizes and training data amounts. For the 3.3B model trained on 100B tokens, Meta-rater achieves an average score of 54.71, surpassing the random sampling baseline of 52.98 by 1.73. The improvement is particularly pronounced in General Knowledge tasks (67.51 vs 64.22). For the 7.2B model trained on 150B tokens, the gap widens further, with Meta-rater outperforming random sampling by 3.12 (55.24 vs 52.12). These results demonstrate that Meta-rater’s benefits scale effectively to larger models and datasets, consistently delivering more efficient training across different model capacities.
Model Method G.K. C.R. R.C. Avg.
3.3B Random 64.22 53.55 35.28 52.98
Meta-rater 67.51 54.35 36.06 54.71
7.2B Random 65.10 52.01 35.87 52.12
Meta-rater 67.97 54.58 37.14 55.24
Table 3: Performance of 3.3B and 7.2B models with random sampling and Meta-rater on downstream tasks. Abbreviations: G.K. = General Knowledge, C.R. = Commonsense Reasoning, R.C. = Reading Comprehension.
@@ 5 Analysis
Table 4: Downstream task results comparison of naive rater combination methods.
5.1 Effect of Proxy Models
Number.
@@ 5.1 Effect of Proxy Models
Figure 2: Average downstream task performance of Meta-rater with different numbers of proxy models.
Proxy Model Architecture.
Table 11 provides details on our original proxy model architecture. To examine if proxy model architecture affects data selection outcomes, we expand the model size from 18M to 46M parameters by increasing hidden dimensions (256$\to$512) and layer count (2$\to$4), then generate a new set of data score weights. Comparing the datasets selected using these new weights with those from the original proxy model revealed a 94.6% overlap between them. This substantial agreement indicates that Meta-rater’s data selection recommendations remain consistent despite moderate changes to the proxy model size.
@@ 5.2 Effect of Combining Strategies for Quality Scores
We explore two straightforward methods for combining quality scores as alternatives to the Meta-rater: Mean: This approach uses the arithmetic mean of all quality scores, giving equal weight to each score, and Intersection [^2] Due to strict selection criteria, sufficient data for pre-training could only be obtained with 4 QuRating raters and 4 PRRC raters.: This method selects data that meet the criteria for all quality scores. While both methods provide ways to combine quality scores, they result in only slight performance improvements. The proposed Meta-rater outperforms the Mean method, achieving an average score of 47.01 compared to 44.65 with 25 raters. A similar gap is observed between Meta-rater and the Intersection method with PRRC raters (46.35 vs 45.17).
We believe this superiority stems from key limitations of these simple combinations. The Mean approach assumes equal importance among all raters, which leads to suboptimal results when there are imbalances in scoring. As shown in Appendix G, individual quality score distributions vary significantly, making uniform weighting ineffective. The Intersection method results in excessive data elimination due to strict filtering criteria, where a single low score can exclude data even if other raters provide high scores. In contrast, Meta-rater’s weighted aggregation dynamically adjusts rater contributions while preserving data integrity. The complexity of optimal weight calibration becomes evident when examining the 25-dimensional weight space. We performed PCA to visualize the loss surface in Figure 5, which reveals multiple local minima. This explains why simple approaches like uniformly weighting scores underperform compared to Meta-rater’s learned weights, as the optimal region forms a relatively small "valley" in the weight space.
@@ 6 Conclusion
We present Meta-rater, a multi-dimensional framework that integrates quality metrics to identify optimal pre-training data for LLMs. Our evaluations demonstrate that Meta-rater consistently outperforms existing data selection methods across downstream tasks, with benefits that scale to larger models up to 7.2B parameters. These results confirm that Meta-rater successfully improves both efficiency and performance in LLM pre-training.